The sum of two numbers is 36 and the difference is 6 . What are the numbers?

Answer by checkley77(12844)

X+Y=36
X-Y=6 ADD.
---------------
2X=42
X=42/2
X=21 ANS.
21+Y=36
Y=36-21
Y=15 ANS.
PROOF:
21-15=6
6=6